# Investigating Factuality in Long-Form Text Generation: The Roles of Self-Known and Self-Unknown

### v1

@@ A.1 LLMs
A.2 Prompt Templates
A.2.1 Prompts Self-known and Self-unknown
@@
A.2.2 Prompts for Creating the Question-answer Pair
@@
A.3 Rules for Filtering Generations
@@
A.4 Automatic Tool
@@
A.5 Details on Computing Experimental Result For each LLM
@@
We feed a prompt “Tell me a bio of $<$entity$>$” to the LLM and take the generation. 500 human entities (Min et al., 2023) are used to generate these biographies.
@@
For lots of LLMs, a biography is not provided if they think they do not have enough detailed information to provide a biography. We implement rules to filter out these generations [^10] The useful rules are shown in Section A.3. .
@@ A.6 More results
Document [0] Jessie Mae Brown Beavers Jessie Mae Brown Beavers (March 18, 1923 – September 6, 1989) was an American journalist based in Los Angeles, California. She was an editor at the ”Los Angeles Sentinel” from 1949 to 1989, and served sixteen years on the city’s Human Relations Commission, beginning with her 1973 appointment by mayor Tom Bradley.Early life. Jessie Mae Brown was born in Los Angeles, the daughter of Arnetta Hoyt Brown, a Baptist deaconess. She attended the University of California, Los Angeles, where she earned a bachelor’s degree in sociology.Career. Brown was editor of the family section of the ”California Eagle” from 1944 to 1949, when she joined the staff of the ”Los Angeles Sentinel” as an editor. In 1966 she was one of the organizers and leaders of the Los Angeles chapter of the National Association of Media Women. In 1969, she was given the Outstanding Woman in Journalism Award by the University of Southern California chapter of Theta Sigma Phi. Beavers served on the Los Angeles Human Relations Commission for sixteen years, after she was appointed by mayor Tom Bradley in 1973. On the commission, she worked closely with fellow commissioner Toshiko S. Yoshida during the American bicentennial

### v2

@@ Appendix A LLMs
Appendix B Prompt Templates
B.1 Prompts Self-known and Self-unknown
@@
B.2 Prompts for Creating the Question-answer Pair
@@
Appendix C Rules for Filtering Generations
@@
Appendix D Automatic Tool Results
See Figure 7
@@ Appendix D Automatic Tool Results
Appendix E Details on Computing Experimental Result For each LLM
@@
We feed a prompt “Tell me a bio of <<entity>>” to the LLM and take the generation. 500 human entities (Min et al., 2023) are used to generate these biographies.
@@
For lots of LLMs, a biography is not provided if they think they do not have enough detailed information to provide a biography. We implement rules to filter out these generations [^8] The useful rules are shown in Section C. .
@@ Appendix F More results
Table 1 in the Appendix presents two results for various LLMs: the average number of atomic claims per generation and the filtered rate. The filtered rate represents the percentage of instances where the LLMs do not provide valuable responses, often due to perceiving insufficient information to generate a meaningful answer. We notice that the behavior of Claude-3-opus and Gemini-1.5-pro is more conservative. These models frequently decide not to provide a valuable response, instead stating something like “I do not have enough verified information”.
@@
Document [0] Jessie Mae Brown Beavers Jessie Mae Brown Beavers (March 18, 1923 – September 6, 1989) was an American journalist based in Los Angeles, California. She was an editor at the "Los Angeles Sentinel" from 1949 to 1989, and served sixteen years on the city’s Human Relations Commission, beginning with her 1973 appointment by mayor Tom Bradley.Early life. Jessie Mae Brown was born in Los Angeles, the daughter of Arnetta Hoyt Brown, a Baptist deaconess. She attended the University of California, Los Angeles, where she earned a bachelor’s degree in sociology.Career. Brown was editor of the family section of the "California Eagle" from 1944 to 1949, when she joined the staff of the "Los Angeles Sentinel" as an editor. In 1966 she was one of the organizers and leaders of the Los Angeles chapter of the National Association of Media Women. In 1969, she was given the Outstanding Woman in Journalism Award by the University of Southern California chapter of Theta Sigma Phi. Beavers served on the Los Angeles Human Relations Commission for sixteen years, after she was appointed by mayor Tom Bradley in 1973. On the commission, she worked closely with fellow commissioner Toshiko S. Yoshida during the American bicentennial